solve for x
0.25mg: IML :: 625 mcg : XML 0 100 100 10 al al 000 roll mboo N -
Answer #1

Ans) 0.25 mg = 1 ml

72.5 mcg = x

( 72.5 micrograms =0.0725 milligram)

As 1 mg = 1 mcg/1000

x = 0.0725 ×1 / 0.25 = 725 × 100/ 25 × 10000 = 0.29 ml

x= 0.29 ~ 0.3 ml